# Testing by Betting while Borrowing and Bargaining

### v1

@@ 4 Gross Wealth as Evidence
We first establish conditions on which Statistician’s gross wealth $\{W_{t}\}$, with liabilities undeducted, can act as evidence against $\mathbb{P}_{1/2}$.
@@
The random variable $L_{\tau}$ is integrable under $\mathbb{P}_{1/2}$: $L=\mathbb{E}_{1/2}L_{\tau}<\infty$.
@@
The process $\{L_{t}\}$ is bounded in expectation under $\mathbb{P}_{1/2}$: $L=\sup_{t}\mathbb{E}_{1/2}L_{t}<\infty$.
@@
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.6 and 4.1. Then, $(1+L)^{-1}W_{\tau}$ is an e-value under $\mathbb{P}_{1/2}$, thus the wealth $W_{\tau}$ is a $(1+L,0)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumptions 3.5 and 4.2, the wealth process $\{W_{t}\}$ is a $(1+L,0)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
@@
(i) Note that $W_{\tau}$ is a nonnegative random variable and $\mathbb{E}_{1/2}W_{\tau}=\mathbb{E}_{1/2}N_{\tau}+\mathbb{E}_{1/2}L_{\tau}\leqslant 1+L$. Therefore, by Markov’s inequality, $\mathbb{P}_{1/2}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$.
(ii) Since Assumption 3.5 holds, we learn from Proposition 3.1 that the wealth process $\{W_{t}\}$ is a nonnegative submartingale. Therefore, letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{1/2}W_{\tau}\leqslant\mathbb{E}_{1/2}W_{T}=\mathbb{E}_{1/2}N_{T}+\mathbb{E}_{1/2}L_{T}\leqslant N_{0}+L=1+L$. Therefore by Markov’s inequality, $\mathbb{P}_{1/2}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$, concluding the proof. ∎
Proposition 4.3 states that the total gross wealth that Statistician gathers over the bets counts as evidence against $\mathbb{P}_{1/2}$, discounted by a factor that corresponds to the average liabilities over all possible coin-tosses. It is easy to see that we can slightly weaken the assumptions of Part (ii) of Proposition 4.3, allowing Statistician to pay back debts but placing instead an upper bound on the expected total actual borrowings.
@@
$B=\sup_{t}\mathbb{E}_{1/2}\sum_{i=1}^{t}\beta_{i}^{+}<\infty,$ (11)
the wealth process $\{W_{t}\}$ is a $(1+B,0)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
@@
Consider the increasing process $B_{t}=\sum_{i=1}^{t}\beta_{i}^{+}$ and $V_{t}=N_{t}+B_{t}$. Then, $V_{t}\geqslant W_{t}$ and $\{V_{t}\}$ is a submartingale. Letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{1/2}W_{\tau}\leqslant\mathbb{E}_{1/2}V_{\tau}\leqslant\mathbb{E}_{1/2}V_{T}=\mathbb{E}_{1/2}N_{T}+\mathbb{E}_{1/2}B_{T}\leqslant 1+B$. Markov’s inequality then completes the proof. ∎
@@ 5 Net Wealth as Evidence
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.6 and 5.1. Then, $(N_{\tau}-N_{\min})/(1-N_{\min})$ is an e-value under $\mathbb{P}_{1/2}$, thus the net wealth $N_{\tau}$ is a $(1-N_{\min},N_{\min})$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumption 5.2, the net wealth process $\{N_{t}\}$ is a $(1-N_{\min},N_{\min})$-sequential tail evidence against $\mathbb{P}_{1/2}$.
@@
(i) follows from $\mathbb{E}_{1/2}(N_{\tau}-N_{\min})=1-N_{\min}$, and $N_{\tau}-N_{\min}\geqslant 0$; and (ii) follows from the fact that $\{N_{t}-N_{\min}\}$ is a nonnegative martingale. ∎
@@
(i) Suppose a bounded stopping time $\tau$ satisfies Assumption 5.4. Then, $(\tilde{N}_{\tau}-G)/(1-G)$ is an e-value under $\mathbb{P}_{1/2}$, thus the sub-net wealth $\tilde{N}_{\tau}$ is a $(1-G,G)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$. (ii) Under Assumption 5.5, the sub-net wealth process $\{\tilde{N}_{t}\}$ is a $(1-G,G)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
@@
$\mathbb{E}_{1/2}(W_{t}-\tilde{L}_{t}|\mathcal{F}_{t-1})=\mathbb{E}_{1/2}(W_{t}-L_{t}|\mathcal{F}_{t-1})+(L_{t}-\tilde{L}_{t})=(W_{t-1}-L_{t-1})+(L_{t}-\tilde{L}_{t})$ (13)
@@
Consider a two-round betting with borrowings $\beta_{1}=\beta_{2}=1$ and bets $\lambda_{1}=\lambda_{2}=1/2$. A sub-liabilities $\tilde{L}_{2}=\beta_{1}+(2-X_{1})\beta_{2}$ is defined to penalize borrowing when already at a loss. We can see from Table 1 that it is possible that the e-value corresponding to the sub-net wealth is larger.
@@ 6 Borrowing as Averaging
Indeed, a direct calculation shows that $\mathbb{E}(W_{\tau_{n+1}}-W_{\tau_{n}}|\mathcal{F}_{\tau_{n}})=\beta_{\tau_{n}+1}$. Note that the e-values $E_{n}$ can be understood without reasoning about borrowing, as no borrowing occurs besides $\beta_{\tau_{n}+1}$, which is treated here as an initial capital.
@@
each $E_{n}$ satisfying $\mathbb{E}_{1/2}(E_{n}|\mathcal{F}_{\tau_{n}})=1$. Note that each $\beta_{\tau_{n}+1}$ is $\mathcal{F}_{\tau_{n}}$-measurable, a weighted average over (18) can be taken, where each $E_{n}$ receiving weight $\beta_{\tau_{n}+1}$, to obtain another e-value,
@@
$\mathbb{E}_{1/2}(E|\mathcal{F}_{\tau_{B}})=\frac{\mathbb{E}_{1/2}\{\sum_{n=0}^{B-1}\mathbb{E}_{1/2}(\beta_{\tau_{n}+1}E_{n}|\mathcal{F}_{\tau_{n}})|\mathcal{F}_{\tau_{B}}\}}{L}=\frac{\sum_{n=0}^{B-1}\mathbb{E}_{1/2}(\beta_{\tau_{n}+1}|\mathcal{F}_{\tau_{B}})}{L}=1.$ (20)
@@
We thus conclude that $E=E^{\prime}$, i.e., the two different perspectives yield the same e-value.
@@ 7 The Futile Borrower
$\mathcal{E}(Y)=\sup_{a,b}\left\{\mathbb{E}_{\mathbb{Q}}(aY+b):\mathbb{E}_{\mathbb{P}}(aY+b\geqslant x)\leqslant x^{-1}\right\},$ (26)
@@
$\mathcal{E}(Y)=\sup_{a,b}\left\{a\mathbb{E}_{\mathbb{Q}}(Y)+b:\mathbb{E}_{\mathbb{P}}(aY+b\geqslant x)\leqslant x^{-1}\right\}.$ (27)
@@
$\mathcal{E}((1+\beta)Y-\beta)=\sup_{a,b}\left\{a(1+\beta)\mathbb{E}_{\mathbb{Q}}(Y)+b-a\beta:\mathbb{E}_{\mathbb{P}}(a(1+\beta)Y+b-a\beta\geqslant x)\leqslant x^{-1}\right\}.$ (28)
@@ 8 Almost Supermartingales and Bargained Bets
Still working with a general scenario, we equip the measurable space $(\Omega,\mathcal{A})$ with a fixed probability $\mathbb{P}$ and an arbitrary filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ in this section.
Robbins and Siegmund [1971] define the following class of stochastic processes.
@@
$\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})=Z_{t-1}(1+b_{t})+\xi_{t}-\zeta_{t}$ (29)
@@
One can see that any integrable stochastic process $\{Z_{t}\}$ adapted to $\{\mathcal{F}_{t}\}$ is an almost supermartingale with respect to some appropriately chosen triplet of predictable processes. For example, $b_{t}=0$, $\xi_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{+}$, and $\zeta_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{-}$. The following maximal inequality holds for nonnegative almost supermartingales.
@@
All these examples, however, are special cases of Proposition 8.2 with the $\{b_{t}\}$ terms set to zero. When could the $\{b_{t}\}$ be non-zero? Let us consider the scenario where Casino misprices the bet – either by mistake or on purpose, for example, when Statistician has successfully bargained for a higher return per dollar at stake.
In all our previous exposition, the wealth of Statistician evolves as
$W_{t}=(W_{t-1}+\beta_{t})\cdot(1+\lambda_{t}X_{t}),$ (31)
because Casino offers double-or-nothing bets, on both heads and tails, in which case the bets are fairly priced to the effect that no arbitrage opportunities exist for Statistician to win risk-free money. Let us instead consider the bets offered on $X_{t}$ being $(2+2b_{t})$-times-or-nothing for some $b_{t}>0$: for any dollar placed on either heads or tails333 Of note, these mispriced bets we consider here are offered two-sided in symmetry. If only one side is mispriced, or if two sides are priced differently (which normally would fairly price a non-$\mathbb{P}_{1/2}$ null), the condition for the existence of arbitrage shall be different accordingly., Casino will pay back $2+2b_{t}$ dollars if correctly guessed. In this case, we note, Statistician is able to arbitrage by putting $1/2$ on heads and $1/2$ on tails, increasing wealth from 1 to $1+b_{t}$ regardless of the actual distribution $\mathbb{P}_{p}$ and the actual outcome of the coin toss $X_{t}$.
Statistician’s wealth in this mispriced case evolves as
$W_{t}=(W_{t-1}+\beta_{t})\cdot(1+b_{t})(1+\lambda_{t}X_{t}),$ (32)
where $\lambda_{t}\in[-1,1]$ as before. The multiplier $(1+b_{t})(1+\lambda_{t}X_{t})$ is interpreted as follows: Statistician allocates $|\lambda_{t}|$ portion of their wealth into risky bets on either heads ($\lambda_{t}>0$) or tails ($\lambda_{t}<0$), paying off $(1+b_{t})|\lambda_{t}|+(1+b_{t})\lambda_{t}X_{t}$ as per Casino’s protocol; and $1-|\lambda_{t}|$ of their wealth on the risk-free 50-50 combination of bets on heads and tails, paying off $(1-|\lambda_{t}|)(1+b_{t})$, as it does not make sense for Statistician to keep this idle cash. In total, therefore, the wealth is to be multiplied by $(1+b_{t})(1+\lambda_{t}X_{t})$.
@@
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b}+c,$ (33)
we say that it is an $(a,b,c)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
@@
$B=\sup_{t}\mathbb{E}_{1/2}\sum_{i=1}^{t}(1+b_{i})\beta_{i}^{+}<\infty,\quad C=\sum_{t=1}^{\infty}b_{t}<\infty,$ (34)
the wealth process $\{W_{t}\}$ is a $(1+B,0,C)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
As can be already seen from the new definition of $(a,b,c)$-sequential tail evidence in the statement, the evidence that the wealth process $\{W_{t}\}$ generates in this case is of lower quality; there is always a $C$ probability no matter how large the gross wealth has been accrued. That is, as long as $C\geqslant 0.05$, the corresponding $0.05$-level sequential test is trivial. We thus seek to develop evidence with net wealth instead, as we have done in Section 5. However, if we still define the net wealth as $W_{t}-(\beta_{1}+\dots+\beta_{t})$, immediately we run into trouble: since arbitrage opportunities exist, Statistician can borrow an unbounded amount of money, put it all in the risk-free 50-50 combination of bets on heads and tails, and take a net gain of $b$ multiplied by the borrowed amount. Not only the wealth but the net wealth increases unbounded and risk-free, with absolutely no insight gained into the verity of $\mathbb{P}_{1/2}$ as the growth is completely independent of the outcomes. The crux of the matter lies in our earlier assumption that Statistician can borrow without interest. When a risk-free instrument exists, financial literature tells us that the interest of borrowing shall be set the same as the risk-free return rate to prevent pathological possibilities in the system.
@@
is the Doob decomposition of $\{W_{t}^{\prime}\}$ into the martingale $\{N_{t}^{\prime}\}$ and the predictable process $\{L_{t}^{\prime}\}$.

### v2

@@ 4 Gross Wealth as Evidence
We first establish conditions on which Statistician’s gross wealth $\{W_{t}\}$, with liabilities undeducted, can act as evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
The random variable $L_{\tau}$ is integrable under $\mathbb{P}_{\mathsf{S}}$: $L=\mathbb{E}_{\mathsf{S}}L_{\tau}<\infty$.
@@
The process $\{L_{t}\}$ is bounded in expectation under $\mathbb{P}_{\mathsf{S}}$: $L=\sup_{t}\mathbb{E}_{\mathsf{S}}L_{t}<\infty$.
@@
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.5 and 4.1. Then, $(1+L)^{-1}W_{\tau}$ is an e-value under $\mathbb{P}_{\mathsf{S}}$, thus the wealth $W_{\tau}$ is a $(1+L,0)$-tail evidence against $\mathbb{P}_{\mathsf{S}}$ at $\tau$. (ii) Under Assumptions 3.4 and 4.2, the wealth process $\{W_{t}\}$ is a $(1+L,0)$-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
(i) Note that $W_{\tau}$ is a nonnegative random variable and $\mathbb{E}_{\mathsf{S}}W_{\tau}=\mathbb{E}_{\mathsf{S}}N_{\tau}+\mathbb{E}_{\mathsf{S}}L_{\tau}\leqslant 1+L$. Therefore, by Markov’s inequality, $\mathbb{P}_{\mathsf{S}}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$.
(ii) Since Assumption 3.4 holds, we learn from Proposition 2.4 that the wealth process $\{W_{t}\}$ is a nonnegative submartingale. Therefore, letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{\mathsf{S}}W_{\tau}\leqslant\mathbb{E}_{\mathsf{S}}W_{T}=\mathbb{E}_{\mathsf{S}}N_{T}+\mathbb{E}_{\mathsf{S}}L_{T}\leqslant N_{0}+L=1+L$. Therefore by Markov’s inequality, $\mathbb{P}_{\mathsf{S}}(W_{\tau}\geqslant x)\leqslant x^{-1}(1+L)$, concluding the proof. ∎
Proposition 4.3 states that the total gross wealth that Statistician gathers over the bets counts as evidence against $\mathbb{P}_{\mathsf{S}}$, discounted by a factor that corresponds to the average liabilities over all possible outcomes of the bets. It is easy to see that we can slightly weaken the assumptions of Part (ii) of Proposition 4.3, allowing Statistician to pay back debts but placing instead an upper bound on the expected total actual borrowings.
@@
$B=\sup_{t}\mathbb{E}_{\mathsf{S}}\sum_{i=1}^{t}\beta_{i}^{+}<\infty,$ (13)
the wealth process $\{W_{t}\}$ is a $(1+B,0)$-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
Consider the increasing process $B_{t}=\sum_{i=1}^{t}\beta_{i}^{+}$ and $V_{t}=N_{t}+B_{t}$. Then, $V_{t}\geqslant W_{t}$ and $\{V_{t}\}$ is a submartingale. Letting $\tau\leqslant T$ be any bounded stopping time, the optional stopping theorem implies that $\mathbb{E}_{\mathsf{S}}W_{\tau}\leqslant\mathbb{E}_{\mathsf{S}}V_{\tau}\leqslant\mathbb{E}_{\mathsf{S}}V_{T}=\mathbb{E}_{\mathsf{S}}N_{T}+\mathbb{E}_{\mathsf{S}}B_{T}\leqslant 1+B$. Markov’s inequality then completes the proof. ∎
@@ 5 Net Wealth as Evidence
(i) Suppose a stopping time $\tau$ satisfies both Assumptions 3.5 and 5.1. Then, $(N_{\tau}-N_{\min})/(1-N_{\min})$ is an e-value under $\mathbb{P}_{\mathsf{S}}$, thus the net wealth $N_{\tau}$ is a $(1-N_{\min},N_{\min})$-tail evidence against $\mathbb{P}_{\mathsf{S}}$ at $\tau$. (ii) Under Assumption 5.2, the net wealth process $\{N_{t}\}$ is a $(1-N_{\min},N_{\min})$-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
(i) follows from $\mathbb{E}_{\mathsf{S}}(N_{\tau}-N_{\min})\leqslant 1-N_{\min}$, and $N_{\tau}-N_{\min}\geqslant 0$; and (ii) follows from the fact that $\{N_{t}-N_{\min}\}$ is a nonnegative supermartingale. ∎
@@
(i) Suppose a bounded stopping time $\tau$ satisfies Assumption 5.4. Then, $(\tilde{N}_{\tau}-G)/(1-G)$ is an e-value under $\mathbb{P}_{\mathsf{S}}$, thus the sub-net wealth $\tilde{N}_{\tau}$ is a (1G,G)(1-G,G)-tail evidence against $\mathbb{P}_{\mathsf{S}}$ at $\tau$. (ii) Under Assumption 5.5, the sub-net wealth process $\{\tilde{N}_{t}\}$ is a (1G,G)(1-G,G)-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
$\mathbb{E}_{\mathsf{S}}(W_{t}-\tilde{L}_{t}|\mathcal{F}_{t-1})=\mathbb{E}_{\mathsf{S}}(W_{t}-L_{t}|\mathcal{F}_{t-1})+(L_{t}-\tilde{L}_{t})=(W_{t-1}-L_{t-1})+(L_{t}-\tilde{L}_{t})$ (15)
@@
Consider a two-round betting on coin tosses $X_{1},X_{2}\in\{-1,1\}$ with borrowings $\beta_{1}=\beta_{2}=1$ and bet portions $\lambda_{1}=\lambda_{2}=1/2$. A sub-liabilities $\tilde{L}_{2}=\beta_{1}+(2-X_{1})\beta_{2}$ is defined to penalize borrowing when already at a loss. We can see from Table 1 that it is possible that the e-value corresponding to the sub-net wealth is larger.
@@
6 Current Evidence
In previous sections, we have constructed e-values based on assumptions that Statistician knows either an upper bound of $\mathbb{E}_{\mathsf{S}}L_{t}$ or an almost sure lower bound of $N_{t}$. Final wealth is adjusted by these bounds into e-values. These assumptions, however, both require knowing some information regarding how much the borrowing situation would be in hypothetical worlds. Naturally, we would want evidence against $\mathbb{P}_{\mathsf{S}}$ based purely on what has actually happened in the current world. We call e-values and supermartingales that require no knowledge of other worlds “current” e-values and supermartingales. We present the following example of an assumption-free current supermartingale, then discuss its implications. The supermartingale requires no assumptions other than the gross wealth $W_{t}$ always stays positive, i.e. non-zero; it is adjusted from the gross wealth $W_{t}$ by a purely empirical quantity. The positivity of $W_{t}$, we note, is a non-material assumption and is equivalent to disallowing “all-in bets” ($B_{t}>0$) and “paying back debts with all wealth” ($\beta_{t}>-W_{t-1}$).
Theorem 6.1.
Suppose $W_{t}>0$ for all $t$. We define the “instantaneuous leverage ratio” before initiating the bet $B_{t}$ as $\rho_{t}=(W_{t-1}+\beta_{t})/W_{t-1}$, which is positive and $\mathcal{F}_{t-1}$ measurable. The process
$V_{t}=\frac{W_{t}}{\rho_{1}\dots\rho_{t}}$ (17)
is a nonnegative supermartingale under $\mathbb{P}_{\mathsf{S}}$ and martingale under $\mathbb{P}_{\mathsf{M}}$.
Proof.
Since $\rho_{1},\dots,\rho_{t}$ are all $\mathcal{F}_{t-1}$-measurable,
$\mathbb{E}_{\mathsf{M}}\left(\frac{W_{t}}{\rho_{1}\dots\rho_{t}}\middle|\mathcal{F}_{t-1}\right)=\frac{\mathbb{E}_{\mathsf{M}}(W_{t}|\mathcal{F}_{t-1})}{\rho_{1}\dots\rho_{t}}\stackrel{{\scriptstyle(*)}}{{=}}\frac{W_{t-1}+\beta_{t}}{\rho_{1}\dots\rho_{t}}=\frac{W_{t-1}\rho_{t}}{\rho_{1}\dots\rho_{t}}=\frac{W_{t-1}}{\rho_{1}\dots\rho_{t-1}},$ (18)
and the equality ()(*) becomes $\leqslant$ if under $\mathbb{E}_{\mathsf{S}}$. This concludes the proof. ∎
It is worth reiterating that Theorem 6.1 requires no additional assumption beyond our problem setup in Section 2, besides $B_{t}>0$ and $\beta_{t}>-W_{t-1}$ which prevent division by zero: Statistician cannot go all in or pay back all their wealth. A closer look at $V_{t}$, however, reveals that $V_{t}$ is not simply penalizing the accrued wealth $W_{t}$ by the leverage ratios $\rho_{1},\dots,\rho_{t}$, but itself a wealth martingale corresponding to another betting/borrowing strategy: $\{V_{t}\}$ is what the wealth would have been if Statistician never borrowed, but used the same bets $\{B_{t}\}$ as they actually do (e.g. same betting strategy $\{\lambda_{t}\}$ when coin betting). Mathematically,
$W_{0}=1,\quad W_{t}=(W_{t-1}+\beta_{t})\cdot B_{t};$ (19)
$V_{0}=1,\quad V_{t}=(V_{t-1}+0)\cdot B_{t}.$ (20)
@@ 7 Borrowing as Averaging
Indeed, a direct calculation shows that $\mathbb{E}_{\mathsf{M}}(W_{\tau_{n+1}}-W_{\tau_{n}}|\mathcal{F}_{\tau_{n}})=\beta_{\tau_{n}+1}$. Note that the e-values $E_{n}$ can be understood without reasoning about borrowing, as no borrowing occurs besides $\beta_{\tau_{n}+1}$, which is treated here as an initial capital.
@@
each $E_{n}$ satisfying $\mathbb{E}_{\mathsf{M}}(E_{n}|\mathcal{F}_{\tau_{n}})=1$. Note that each $\beta_{\tau_{n}+1}$ is $\mathcal{F}_{\tau_{n}}$-measurable, a weighted average over (24) can be taken, where each $E_{n}$ receiving weight $\beta_{\tau_{n}+1}$, to obtain another e-value,
@@
$\mathbb{E}_{\mathsf{M}}(E|\mathcal{F}_{\tau_{B}})=\frac{\mathbb{E}_{\mathsf{M}}\{\sum_{n=0}^{B-1}\mathbb{E}_{\mathsf{M}}(\beta_{\tau_{n}+1}E_{n}|\mathcal{F}_{\tau_{n}})|\mathcal{F}_{\tau_{B}}\}}{L}=\frac{\sum_{n=0}^{B-1}\mathbb{E}_{\mathsf{M}}(\beta_{\tau_{n}+1}|\mathcal{F}_{\tau_{B}})}{L}=1.$ (26)
@@
We thus conclude that $E=E^{\prime}$, i.e., the two different perspectives yield the same e-value. The case under $\mathbb{P}_{\mathsf{S}}$ is analogous, replacing a few equalities above with inequalities.
@@ 8 The Futile Borrower
$\mathcal{E}(Y)=\sup_{a,b}\left\{\mathbb{E}_{\mathbb{Q}}(aY+b):\mathbb{P}(aY+b\geqslant x)\leqslant x^{-1}\right\},$ (32)
@@
$\mathcal{E}(Y)=\sup_{a,b}\left\{a\mathbb{E}_{\mathbb{Q}}(Y)+b:\mathbb{P}(aY+b\geqslant x)\leqslant x^{-1}\right\}.$ (33)
@@
$\mathcal{E}((1+\beta)Y-\beta)=\sup_{a,b}\left\{a(1+\beta)\mathbb{E}_{\mathbb{Q}}(Y)+b-a\beta:\mathbb{P}(a(1+\beta)Y+b-a\beta\geqslant x)\leqslant x^{-1}\right\}.$ (34)
@@ 9 Almost Supermartingales and Bargained Bets
We now discuss a more generalized setup of borrowing and betting. To set the stage, we note that some of the earlier results presented in this paper can be derived alternatively via the concept of “almost supermartingales”, which Robbins and Siegmund [1971] define as the following class of stochastic processes.
@@
$\mathbb{E}(Z_{t}|\mathcal{F}_{t-1}){\leqslant}Z_{t-1}(1+b_{t})+\xi_{t}-\zeta_{t}$ (35)
@@
One can see that any integrable stochastic process $\{Z_{t}\}$ adapted to $\{\mathcal{F}_{t}\}$ is an almost supermartingale, with respect to some appropriately chosen triplet of predictable processes. For example, $b_{t}=0$, $\xi_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{+}$, and $\zeta_{t}=(\mathbb{E}(Z_{t}|\mathcal{F}_{t-1})-Z_{t-1})^{-}$, and (35) holds with with equality. The following maximal inequality holds for nonnegative almost supermartingales.
@@
All these examples, however, are special cases of Proposition 9.2 with the $\{b_{t}\}$ terms set to zero. When could the $\{b_{t}\}$ be non-zero? Let us first write down the formal definition as a generalization of Definition 2.3, and then discuss its meaning.
Definition 9.3 (Borrowed and Bargained NSM Betting).
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space with filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. Let $\{B_{t}\}_{t\geqslant 1}$, $\{b_{t}\}_{t\geqslant 1}$ be nonnegative processes adapted and predictable w.r.t. $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ respectively, and $\{\beta_{t}\}_{t\geqslant 1}$ a process predictable w.r.t. $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ such that $\beta_{t}\geqslant-W_{t-1}$ for each $t$. The wealth process $\{W_{t}\}_{t\geqslant 0}$ is defined as $W_{0}=1$ and
$W_{t}=(W_{t-1}+\beta_{t})\cdot(1+b_{t})B_{t}.$ (37)
The supermartingale null $\mathbb{P}_{\mathsf{S}}$ under which $\mathbb{E}_{\mathsf{S}}(B_{t}|\mathcal{F}_{t-1})\leqslant 1$ for all $t\geqslant 1$, and the martingale null $\mathbb{P}_{\mathsf{M}}$ under which $\mathbb{E}_{\mathsf{M}}(B_{t}|\mathcal{F}_{t-1})=1$ for all $t\geqslant 1$ are as before.
Apparently, Definition 2.3 is the special case of Definition 9.3 with $b_{t}=0$. With $b_{t}>0$, this setup can be interpreted as the scenario where the bet takes place alongside a risk-free instrument; or equivalently, scenario where Statistician has successfully “bargained” for enjoying the same potential payoff by risking less capital. Let us elucidate this by extending on the coin betting examples mentioned in Section 2.
With coin tossing $X_{t}\in\{-1,1\},B_{t}=1+\lambda_{t}X_{t}$, the previous setup is that Casino offers double-or-nothing bets on both heads and tails, in which case the bets are fairly priced to the effect that no arbitrage opportunities exist for Statistician to win risk-free money. Definition 9.3 corresponds to the case where bets offered on $X_{t}$ are $(2+2b_{t})$-times-or-nothing: for any dollar placed on either heads or tails, Casino will pay back $2+2b_{t}$ dollars if Statistician correctly guessed the side. In this case, we note, Statistician is able to arbitrage by putting $1/2$ on heads and $1/2$ on tails, increasing wealth from 1 to $1+b_{t}$ regardless of the actual distribution $\mathbb{P}$ and regardless of the actual outcome of the coin toss $X_{t}$. With Definition 9.3, Statistician’s strategy at time t1t-1 is as follows: Statistician allocates $|\lambda_{t}|$ portion of their post-borrow wealth $W_{t-1}+\beta_{t}$ into risky bets on heads ($\lambda_{t}>0$) or tails ($\lambda_{t}<0$), paying off
$(W_{t-1}+\beta_{t})\cdot|\lambda_{t}|(2+2b_{t})\mathbbmss{1}_{\{\lambda_{t}X_{t}>0\}}$ (38)
as per Casino’s protocol; and $1-|\lambda_{t}|$ of their post-borrow wealth $W_{t-1}+\beta_{t}$ on the risk-free 50-50 combination of bets on heads and tails, paying off
$(W_{t-1}+\beta_{t})\cdot(1-|\lambda_{t}|)(1+b_{t})$ (39)
as it does not make sense for Statistician to keep this idle cash uninvested. In total, therefore, the wealth after the payoff is
$(W_{t-1}+\beta_{t})\cdot(1+b_{t})\{2|\lambda_{t}|\mathbbmss{1}_{\{\lambda_{t}X_{t}>0\}}+(1-|\lambda_{t}|)\}=(W_{t-1}+\beta_{t})\cdot(1+b_{t})(1+\lambda_{t}X_{t}),$ (40)
agreeing with Definition 9.3.
The general setup with bet process $\{B_{t}\}$ admits a simple economic interpretation: cash earns interest at rate $b_{t}$. Here, $B_{t}$ is no longer what 1 unit of wealth invested in Statistician’s portfolio at time t1t-1 becomes at time $t$; but the (t1)(t-1)-present value of what 1 unit of wealth invested in the portfolio becomes at time $t$, discounted by $1+b_{t}$. Statistician’s wealth thus becomes $W_{t-1}\cdot B_{t}\cdot(1+b_{t})$ at time $t$, coinciding with Definition 9.3.
@@
$\mathbb{P}_{\mathsf{S}}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b}+c,$ (41)
we say that it is an $(a,b,c)$-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
@@
$B=\sup_{t}\mathbb{E}_{\mathsf{S}}\sum_{i=1}^{t}(1+b_{i})\beta_{i}^{+}<\infty,\quad C=\sum_{t=1}^{\infty}b_{t}<\infty,$ (42)
the wealth process $\{W_{t}\}$ is a $(1+B,0,C)$-sequential tail evidence against $\mathbb{P}_{\mathsf{S}}$.
As can already be seen from the new definition of $(a,b,c)$-sequential tail evidence in the statement, the evidence that the wealth process $\{W_{t}\}$ generates in this case is of lower quality; there is always a $C$ probability no matter how large the gross wealth has been accrued. That is, as long as $C\geqslant 0.05$, the corresponding $0.05$-level sequential test is trivial. We thus seek to develop evidence with net wealth instead, as we have done in Section 5. However, if we still define the net wealth as $W_{t}-(\beta_{1}+\dots+\beta_{t})$, immediately we run into trouble: since arbitrage opportunities exist, Statistician can borrow an unbounded amount of money, put it all in the risk-free instrument e.g. the 50-50 combination of bets on heads and tails, and take a net gain of $b$ multiplied by the borrowed amount. Not only the wealth but the net wealth increases unbounded and risk-free, with absolutely no insight gained into the verity of $\mathbb{P}_{\mathsf{S}}$ as the growth is completely independent of the outcomes. The crux of the matter lies in our earlier assumption that Statistician can borrow without interest. When a risk-free instrument exists, the interest of borrowing shall be set at the same rate as the risk-free return rate to prevent pathological possibilities in the system.
@@
where $\{N_{t}^{\prime}\}$ is a supermartingale under $\mathbb{P}_{\mathsf{S}}$ and $\{L_{t}^{\prime}\}$ is predictable. Further, under $\mathbb{P}_{\mathsf{M}}$, $\{N_{t}^{\prime}\}$ is a martingale and (44) is the Doob decomposition of the process $\{W_{t}^{\prime}\}$.